OpenStudy (victore2usir):

A mixture problem. A chemistry experiment calls for a 50% solution of copper sulfate. Sarah has 60 milliliters of 30% solution. How many milliliters of 60% solution should she add to obtain the required 50% solution? I found this question online and I forgot the steps to solving such a problem, could someone help me solve this question step-by-step?

OpenStudy (tkhunny):

Equate something. Start: 60 ml 30% CUSO4 (How much copper sulfate?) Add: x ml 60% (How much copper sulfate?) Stop: (60+x) ml 50% (How much copper sulfate.)

OpenStudy (dumbcow):

To expand on last post. When something is given as a percent solution, you can always multiply the percentage by the total amount of liquid to get the amount of the concentrated solution (in this case copper sulfate) start 60* 30% = 18 ml of copper sulfate add x*60% = 0.6x ml copper sulfate finish (60+x) * 50% = 0.5x + 30 Set up equation using only the concentrated amounts ---> 18+0.6x = 0.5x+30

OpenStudy (victore2usir):

I'll try that, and thanks for the reply!